OpenStudy (anonymous):

write the equation with a slope of -2/3 through the piont(4,-5) in simplest form.

OpenStudy (anonymous):

\[m=\frac{y-y _{1}}{x-x _{1}}\]

OpenStudy (anonymous):

-2/3=y+5/(x-4)

OpenStudy (anonymous):

y=-2/3x-7/3

OpenStudy (anonymous):

but in y=mx+b form..

OpenStudy (mertsj):

That is y = mx+b form

OpenStudy (mertsj):